Grade 12th passInorganic Chemistry

what is the oxidation number of Fe in K3[Fe(C2O4)3]?
pls explain me step wise step......i am didnt get answer

Pareena Yadav
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago

K3[Fe(C2O4)4] is the given coordination compound.

The overall charge is 0.

Let the oxidation number be x.

The charge in potassium is +1.

The charge on C2O4 (ox) is -2.

3*1+1*x+(-2)*3=0

x=6-3=3. Hence, the oxidation number of Fe is +3